OpenStudy (anonymous):

What is the diameter of a circle with a circumference of 14π? thanks

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

the formula for the circumference is pie times diameter.

OpenStudy (unklerhaukus):

\[C=πd\] \[14π=πd\] \[ d=14\]

OpenStudy (anonymous):

\[\pi D= 14 \pi\]\[\pi D / \pi = 14 \pi / \pi\]\[D = 14\]

OpenStudy (anonymous):

thanks so very much for the help

OpenStudy (anonymous):

:)